# Imitation Learning for Combinatorial Optimisation under Uncertainty

### v1

@@ 1 Introduction
Machine learning (ML) has recently emerged as a powerful complementary tool for tackling CO problems, particularly when traditional techniques scale poorly. In high-dimensional or uncertain environments, ML methods can learn approximate decision rules by exploiting patterns in data, enabling faster and more adaptive decision making. This capability is especially valuable in real-time operational settings. Reinforcement learning (RL) [sutton1998reinforcement], in particular, has demonstrated strong potential by learning decision policies through interaction with an environment, using reward feedback to iteratively improve performance. Such approaches are well-suited to settings where explicit models are unavailable, incomplete, or prohibitively complex.
Alongside RL, Imitation learning (IL) [hussein2017imitation] has gained attention as an attractive alternative, especially in scenarios where defining a suitable reward function is difficult but expert demonstrations are available. Instead of learning through trial and error, IL aims to learn a direct mapping from system states to actions by mimicking expert behaviour. A central challenge in this paradigm is the acquisition of high-quality expert demonstrations, which are often expensive or infeasible to generate due to the computational cost of solving CO problems optimally. This has motivated growing interest in approximate, surrogate, or interactive expert models that can provide informative guidance while remaining computationally tractable. The design and utilisation of such experts remain key open questions at the intersection of ML and combinatorial optimisation.
@@
The concept of imitating an expert policy within CO is relatively recent, and existing work proposes a wide range of approaches for defining or approximating experts. These include offline optimal solutions, simplified surrogate models, and structured or repeated interactions with problem-specific oracles. However, to the best of our knowledge, the literature lacks a systematic understanding of what constitutes an expert and how expert choices influence learning outcomes in IL for CO. To address this gap, this paper introduces a taxonomy of expert design for IL in CO. The taxonomy categorises experts along key dimensions, providing a conceptual framework to guide the principled selection and design of experts in different problem settings.
To support this taxonomy, we present a survey of recent literature that illustrates the diversity of expert construction and usage strategies. In the second part of the paper, we propose a two-stage learning pipeline in which a CO-based expert layer generates demonstrations that are subsequently used to train an ML-based decision model. By systematically varying expert properties within this framework, we conduct a comparative analysis that highlights the relative strengths and limitations of different expert types in the context of a physician scheduling problem. The results provide practical insights into how expert design choices affect learning performance, computational cost, and solution quality in sequential decision-making problems.
@@
The remainder of the paper is organised as follows. Section 2 reviews key concepts in ML and CO for sequential decision making. Section 3 introduces the fundamentals of IL. Section 4 presents the proposed taxonomy of expert policies and situates it within the existing literature. Section 5 describes the problem setting considered in this study, and Section 6 details the proposed methodology for applying IL to sequential decision making under uncertainty. Section 7 presents the experimental results and analysis. Section 8 concludes the paper.
@@ 2 Background
Many real-world combinatorial optimisation problems can be naturally formulated as sequential decision problems (SDPs) [powell2007approximate]. At each decision epoch, an agent selects an action based on the current system state and the information available at that time, while accounting for the impact of the decision on both immediate outcomes and future system evolution. The decision process is inherently dynamic, as new information becomes available over time and the system evolves in response to previously taken actions. The objective is typically to design policies that minimise cumulative cost or maximise cumulative reward over a finite or infinite planning horizon. This modelling framework is particularly well suited to applications that involve uncertainty, temporal structure, and adaptive decision-making.
Sequential decision problems are commonly modelled using the Markov decision process (MDP) framework. We begin by introducing this framework in general terms, outlining its key components and underlying assumptions. We then briefly review exact solution methods, including dynamic programming, and discuss their primary limitation, namely, computational intractability in large-scale problems due to the curse of dimensionality. To support the developments in later sections, we also introduce fundamental concepts from machine learning that are relevant to sequential decision making. Finally, we discuss how machine learning techniques have been employed to develop approximate solution methods for MDPs, offering scalable alternatives when exact approaches are computationally infeasible.
@@ 2.1 Markov Decision Process
A Markov decision process is a mathematical framework used to model decision-making problems in which outcomes are influenced by both random factors and controllable actions [powell2022reinforcement]. The core elements of an MDP are as follows.
@@
  • Decision epoch ($k$) - A decision epoch is a point in time at which the decision maker observes the current system state and selects an action. Let $K$ denote the final decision epoch.
@@
  • State ($x_{k}$) - The state provides a complete description of the system at decision epoch $k$ that is sufficient for making future decisions. The collection of all possible states defines the state space $\mathcal{S}$. When the epoch index is not required, we use $x$ to denote a generic state.
@@
  • Decision variables ($a_{k}$) - A decision variable represents the action taken at epoch $k$. It may be binary, discrete, continuous, or vector-valued, and is subject to feasibility constraints of the form $a_{k}\in A_{k}$.
  • Exogenous information ($\xi_{k}$) - Exogenous information represents uncertainty revealed after a decision is made. This information is not known at the time of decision-making and originates from outside the system being modelled.
@@
  • Cost/Reward Function ($C_{k}(x_{k},a_{k})$) – This function specifies the cost incurred or reward obtained when action $a_{k}$ is taken in state $x_{k}$ at decision epoch $k$.
  • Transition function ($X^{M}(x_{k},a_{k},\xi_{k})$) - The transition function specifies how the system evolves from one state to the next. It defines the equations used to update each component of the state and captures all system dynamics, including the updating of estimates and beliefs.
  • Policy ($\pi$) - A policy is a mapping from a state $x_{k}$ to an admissible action $a_{k}$. Policies may be represented in tabular form or as parameterised functions, for example, using parameters $\theta$. The optimal policy $\pi^{*}$ is defined as
    $=\operatorname*{arg\,min}_{\pi}V^{\pi}(x_{k}).$ (2)
  • Value function - The value function represents the cumulative cost or reward accrued over the planning horizon. Assuming a cost-minimisation objective, the value function from an initial state $x_{0}$ is given by
    $V^{\pi^{*}}(x_{0})=\min_{\pi}\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},\pi(x_{k}))\,\big|\,x_{0}\Bigg\},$ (3)
    where $a_{k}=\pi(x_{k})$. For simplicity, discounting of future costs is omitted, with a primary focus on finite-horizon problems. In the MDP literature, $V^{\pi^{*}}(x_{k})$ is referred to as the optimal value function or optimal cost-to-go from state $x_{k}$ at epoch $k$.
    Given a value function, the optimal policy can be recovered via a one-step greedy minimisation. Value functions for a fixed policy can be computed using dynamic programming methods [puterman2014markov], and Equation˜2 can be solved using value iteration or policy iteration algorithms. Value iteration repeatedly updates state values using the Bellman optimality equation until convergence, after which the optimal policy is derived. Policy iteration alternates between policy evaluation and policy improvement steps. A third approach formulates the MDP as a linear programme, optimising the value function subject to Bellman constraints. While these methods yield exact solutions, they are computationally expensive and often infeasible for large-scale problems due to the three curses of dimensionality [powell2007approximate].
    In some applications, interest lies only in the optimal action at a specific epoch $k$. The optimal action $a_{k}^{*}$ is given by
    $a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\Big\{C_{k}(x_{k},a_{k})+\mathbb{E}[V_{k+1}(x_{k+1})]\,\big|\,x_{k}\Big\},$ (4)
    where $V_{k+1}(x_{k+1})$ denotes the optimal value of the next state.
@@
The nature of a decision-making problem depends on the length of the planning horizon $K$. When $K<\infty$, the problem is finite-horizon, whereas $K=\infty$ corresponds to an infinite-horizon setting. A particularly important special case arises when $K=1$. When uncertainty is present, this setting corresponds to a two-stage stochastic programming problem with a single realisation of uncertainty. In the absence of uncertainty, it reduces to a deterministic optimisation problem. Although such problems are static in nature, they often arise repeatedly in practice. For example, a company may solve an independent delivery routing problem each day, with no direct dependence across days. When $K>1$, the problem becomes a multi-stage or fully sequential decision problem.
@@ 2.2 Machine Learning
Machine learning (ML) is a subfield of artificial intelligence that focuses on designing algorithms capable of learning patterns from data to make predictions or decisions without explicit programming. At a high level, the objective is to infer a function that maps inputs to corresponding outputs. Formally, let $X$ denote the input space, commonly referred to as the feature space, and let $Y$ denote the output space. Each input encodes a collection of descriptive attributes for a data instance, and the task is to produce an appropriate output based on these features. We consider a class of parameterised models represented by functions $f_{\theta}:X\rightarrow Y$, where $\theta\in\Theta$ denotes the set of model parameters. The learning process consists of selecting parameter values that yield an effective input–output mapping, typically guided by empirical data and a chosen performance criterion. Depending on the type of feedback available, learning problems are commonly categorised into three main settings: supervised learning, where labelled data are provided; unsupervised learning, where only unlabelled data are available; and reinforcement learning, where an agent learns through interaction with an environment and receives feedback in the form of rewards.
Supervised learning addresses settings in which the learner is provided with a finite collection of input–output pairs $\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in X$ and $y_{i}\in Y$ denote the input and corresponding output, respectively. The outputs $y_{i}$ are commonly referred to as labels or ground truth. The objective is to approximate the underlying input–output relationship by minimising a loss function $\ell:Y\times Y\rightarrow\mathbb{R}$, which measures the discrepancy between predicted and observed outputs. A widely used example is the mean squared error (MSE), defined as
@@
Since the true data-generating distribution is unknown, the loss is minimised empirically over the observed samples.
Unsupervised learning, by contrast, refers to settings in which only input samples $\{x_{i}\}_{i=1}^{N}$ are available, without associated labels. The goal is to uncover latent structure or meaningful representations within the data. Typical tasks include clustering, dimensionality reduction, and generative modelling. These methods rely on objective functions that promote internal consistency or compactness of the learned representations, such as minimising intra-cluster distances in $k$-means clustering. In the absence of ground-truth labels, evaluation typically relies on internal metrics or qualitative assessments informed by domain knowledge.
Reinforcement learning concerns sequential decision-making problems in which a learning agent interacts with an environment over time. At each time step, the agent observes a state, selects an action according to a policy, and receives feedback in the form of a scalar reward, along with a new state resulting from its action. Unlike supervised learning, explicit target outputs are not provided; instead, the agent must infer desirable behaviour from delayed and often sparse rewards. The objective of reinforcement learning is to learn a policy that maximises the expected cumulative reward over time, typically formalised using frameworks such as MDPs. This setting is particularly well-suited to problems involving control, planning, and adaptation, where actions influence future states and feedback is evaluative rather than instructive.
The learning process involves several interconnected components. First, a model class is selected, such as linear predictors, decision trees, kernel methods, or neural networks, each offering different expressive capabilities. Learning proceeds by specifying an appropriate loss function and employing an optimisation algorithm, such as gradient descent or its variants, to adjust model parameters to minimise the loss on the training data. In addition to model parameters, learning algorithms involve hyperparameters, such as learning rates, regularisation strengths, or architectural choices, which govern the learning dynamics and model complexity. To evaluate generalisation performance and mitigate overfitting, data are typically partitioned into a training set, a validation set for hyperparameter tuning, and a test set for final evaluation. Overfitting occurs when a model fits the training data well but fails to generalise to unseen instances.
@@ 2.3 Machine Learning for MDPs
Machine learning provides a range of approaches for solving MDPs in settings where system dynamics are partially known or entirely unknown. RL and approximate dynamic programming (ADP) are two prominent frameworks in this context. RL is commonly applied in model-free settings where transition dynamics are unknown. Algorithms such as Q-learning and policy gradient methods learn policies directly through interaction with the environment, using sampled trajectories to estimate value functions or policy gradients and to iteratively improve performance.
Approximate dynamic programming, more prevalent in operations research and control, often assumes access to a simulator or generative model of the system. ADP addresses the intractability of exact dynamic programming by approximating value functions or policies using function approximation techniques, including linear basis expansions, regression trees, and neural networks. These approximations enable iterative solution methods such as fitted value iteration, approximate policy iteration, and least-squares temporal difference learning. Although RL and ADP differ in historical origins and terminology, the distinction between the two has become increasingly blurred, particularly with the adoption of deep learning–based methods.
Imitation learning (IL) offers an alternative approach for solving MDPs, especially in situations where the reward function is unknown, difficult to specify, or poorly aligned with the desired behaviour. In this setting, the objective is to learn a policy by leveraging expert-generated state–action trajectories. The simplest form of IL is Behavioural Cloning (BC) [pomerleau1988alvinn], which treats imitation as a supervised learning problem that maps observed states directly to expert actions. However, BC is vulnerable to compounding errors arising from distributional shift, whereby the learner encounters states during deployment that are underrepresented in the training data.
To address this issue, more robust IL methods incorporate expert interaction during learning, allowing the learner to query the expert on states encountered during execution. A related class of methods, known as inverse reinforcement learning (IRL) [ng2000algorithms], seeks to infer a reward function for which the expert policy is optimal. Once such a reward function is recovered, standard reinforcement learning or approximate dynamic programming techniques may be applied to compute a policy.
@@ 3 Imitation Learning: Models and Challenges
Imitation learning (IL) [hussein2017imitation] is a specialised branch of supervised machine learning in which an agent learns a task by mimicking behaviour demonstrated by an expert, to replicate the expert decision-making process from observed examples. Unlike classical supervised learning, where data points are typically assumed to be independent and identically distributed, IL often involves sequential decision making. In such settings, data points are temporally correlated and depend on previously taken actions.
The term expert denotes the source of demonstrations used by the learning agent and may correspond to a human decision maker or an exact algorithm. Demonstrations are commonly provided as trajectories $\tau=\{x_{i},a_{i}^{*}\}_{i=1}^{T}$, where $T$ denotes the trajectory length. An expert dataset can then be defined as $\mathcal{D}=\{\tau_{j}\}_{j=1}^{N}$, where each state $x_{i}$ is generated under the expert policy, that is $x_{i}\sim\pi^{*}$, and the corresponding action is given by $a_{i}^{*}=\pi^{*}(x_{i})$. In a machine learning context, an action is also referred to as a target or label. Standard supervised learning methods can then be applied to $\mathcal{D}$ to train a model that predicts expert-like actions from observed states.
Imitation learning offers a data-driven approach for solving complex SDPs with combinatorial action spaces. A major challenge is the acquisition of high-quality expert actions, as optimal solutions are often computationally expensive. Beyond this, IL must scale to high-dimensional decision spaces and address covariate shift, where the learned policy encounters states during deployment that differ from those seen during training, leading to poor generalisation. This section introduces two IL paradigms, reviews key challenges, and discusses the use of noisy or imperfect experts as alternatives to expensive optimal experts.

### v2

@@ 1 Introduction
Machine learning (ML) has recently emerged as a powerful complementary tool for tackling CO problems, particularly in regimes where traditional exact or heuristic methods scale poorly. In many operational settings, decisions must be made repeatedly and under strict time constraints, making it impractical to solve large optimisation models from scratch at every decision point. In such environments, ML methods can exploit historical data and structural regularities in problem instances to learn approximate decision rules or policies that generalise across scenarios. To illustrate this motivation, consider problems such as assigning incoming patients to physicians, dispatching vehicles to delivery requests, or selecting branching decisions within a mixed-integer solver. Solving a full optimisation model at every decision point can be computationally prohibitive, particularly in real-time applications. Yet, across repeated instances, consistent structural patterns often emerge in how high-quality decisions depend on the system state. Machine learning models can exploit these patterns by learning from historical solutions or expert demonstrations, allowing them to approximate the outcome of the optimisation process efficiently. Once trained, such models can generate near-optimal decisions in milliseconds rather than seconds or minutes, while remaining responsive to evolving system conditions.
Reinforcement learning (RL) [sutton1998reinforcement] has demonstrated strong potential, particularly in learning decision policies through interaction with an environment, utilising reward feedback to improve performance iteratively. Such approaches are well-suited to settings where explicit models are unavailable, incomplete, or prohibitively complex. Alongside RL, Imitation learning (IL) [hussein2017imitation] has gained attention as an attractive alternative, especially in scenarios where defining a suitable reward function is difficult but expert demonstrations are available. Instead of learning through trial and error, IL aims to learn a direct mapping from system states to actions by mimicking expert behaviour. A central challenge in this paradigm is the acquisition of high-quality expert demonstrations, which are often expensive or infeasible to generate due to the computational cost of solving CO problems optimally. This has motivated growing interest in approximate, surrogate, or interactive expert models that can provide informative guidance while remaining computationally tractable. The design and utilisation of such experts remain key open questions at the intersection of ML and combinatorial optimisation.
@@
The idea of imitating expert policies in CO is relatively recent, and existing work employs a wide range of strategies to define or approximate experts. These include using offline optimal solutions, simplified surrogate models, and structured or repeated interactions with problem-specific oracles. We first survey recent literature to illustrate the breadth of approaches to expert construction and utilisation. However, despite this growing body of work, there is limited understanding of what constitutes an expert in IL for CO and how different expert design choices influence learning outcomes.
To address this gap, we introduce a taxonomy of expert design for IL in CO. The proposed taxonomy unifies existing approaches by organising experts along key conceptual dimensions, providing a coherent framework that connects and contextualises a wide range of prior methods. It also serves as a guide for the principled selection and design of experts across different problem settings.
In the second part of the paper, we propose a general two-stage learning pipeline in which a CO-based expert layer generates demonstrations that are then used to train an ML-based decision model. The framework exposes several design choices that correspond directly to the dimensions of the expert taxonomy. We conduct a comparative empirical study using a physician scheduling problem to examine how these choices affect learning performance. The results offer practical insights into the trade-offs between expert quality, computational effort, and solution quality in sequential decision-making problems under uncertainty.
@@
The remainder of the paper is organised as follows. Section 2 reviews key concepts in ML and CO for sequential decision making. Section 3 introduces the fundamentals of IL. Section 4 presents the proposed taxonomy of expert policies and situates it within the existing literature. Section 6 describes the problem setting considered in this study, and Section 5 details the proposed methodology for applying IL to sequential decision making under uncertainty. Section 6.2 presents the experimental results and analysis. Section 8 concludes the paper.
@@ 2 Background
Many real-world CO problems can be naturally formulated as sequential decision problems (SDPs) [powell2007approximate]. At each decision epoch, an agent selects an action based on the current system state and available information, while accounting for both immediate consequences and future system evolution. The process is inherently dynamic, as new information is revealed and the system transitions over time. The objective is to design policies that minimise cumulative cost or maximise cumulative reward over a finite or infinite planning horizon, making SDPs well-suited to problems involving uncertainty, timing, and adaptive decision-making.
@@ 2.1 Markov Decision Process
SDPs are commonly modelled using the Markov decision process (MDP) framework, which provides a formal representation of decision-making under uncertainty [powell2022reinforcement]. An MDP models the interaction between controllable actions and stochastic system dynamics over time. In Table˜1, we summarise the core components of an MDP. Additional details are provided in the appendix.
Table 1: Key components of a Markov decision process
Component Description
@@
Decision epoch ($k$) A time index at which the decision maker observes the system state and selects an action. We focus on finite-horizon problems with final epoch $K<\infty$.
@@
State ($x_{k}$) A complete description of the system at epoch $k$ that is sufficient for future decision making. The set of all states defines the state space $\mathcal{S}$.
@@
Action ($a_{k}$) The decision taken at epoch $k$, which may be discrete, continuous, or vector-valued, and must satisfy feasibility constraints $a_{k}\in\mathcal{A}_{k}$.
Exogenous information ($\xi_{k}$) Random information revealed after action selection, representing uncertainty external to the system.
@@
Cost/Reward function ($C_{k}(x_{k},a_{k})$) The immediate cost incurred or reward received when an action is taken in a given state at epoch $k$.
Transition function ($X^{M}(x_{k},a_{k},\xi_{k})$) The mapping that governs the evolution of the system from one state to the next, capturing the system dynamics.
Policy ($\pi$) A mapping from states to admissible actions. Policies may be deterministic or parameterised functions.
@@
A particularly important special case arises when $K=1$. When uncertainty is present, this setting corresponds to a two-stage stochastic programming problem with a single realisation of uncertainty. In the absence of uncertainty, it reduces to a deterministic optimisation problem. Although such problems are static in nature, they often arise repeatedly in practice. For example, a company may solve an independent delivery routing problem each day, with no direct dependence across days. When $K>1$, the problem becomes a multi-stage or fully SDP.
@@ 2.2 Machine Learning
Machine learning (ML) [hastie2009elements, bishop2006pattern] is a subfield of artificial intelligence concerned with learning patterns from data to make predictions or decisions without explicit programming. At a high level, the objective is to learn a mapping from inputs to outputs. Let $X$ denote the input (feature) space and $Y$ the output space. We consider parameterised models of the form $f_{\theta}:X\rightarrow Y$, where $\theta\in\Theta$ denotes the model parameters. Learning consists of selecting parameters that yield an effective input–output mapping based on observed data and a chosen performance criterion. Depending on the type of feedback available, learning problems are commonly classified into supervised, unsupervised, and reinforcement learning.
Supervised learning considers settings in which the learner is given a finite collection of labelled examples $\{(x_{i},y_{i})\}_{i=1}^{N}$, with $x_{i}\in X$ and $y_{i}\in Y$. The objective is to approximate the underlying input–output relationship by minimising a loss function $\ell:Y\times Y\rightarrow\mathbb{R}$ that measures prediction error. A common example is the mean squared error,
@@
Since the true data-generating distribution is unknown, optimisation is performed empirically over the observed samples.
Unsupervised learning addresses scenarios in which only input samples $\{x_{i}\}_{i=1}^{N}$ are available, without labels. The goal is to uncover latent structure or informative representations of the data. Typical tasks include clustering, dimensionality reduction, and representation learning. Evaluation is generally based on internal metrics or qualitative assessment, as ground truth labels are unavailable.
Reinforcement learning focuses on sequential decision-making problems in which an agent interacts with an environment over time. At each step, the agent observes a state, selects an action, and receives scalar feedback in the form of a reward, together with a new state. Unlike supervised learning, explicit target outputs are not provided; instead, the agent must infer desirable behaviour from delayed and possibly sparse rewards. The objective is to learn a policy that maximises expected cumulative reward, typically formalised using the MDP framework.
The learning process involves several interconnected components. First, a model class is selected, such as linear predictors, decision trees, kernel methods, or neural networks, each offering different expressive capabilities. Learning proceeds by specifying an appropriate loss function and employing an optimisation algorithm, such as gradient descent or its variants, to adjust model parameters to minimise the loss on the training data. In addition to model parameters, learning algorithms involve hyperparameters, such as learning rates, regularisation strengths, or architectural choices, which govern the learning dynamics and model complexity. To evaluate generalisation performance and mitigate overfitting, data are typically partitioned into a training set, a validation set for hyperparameter tuning, and a test set for final evaluation. Overfitting occurs when a model fits the training data well but fails to generalise to unseen instances. We refer the reader to [hastie2009elements] for more information of these elements.
@@ 2.3 Machine Learning for MDPs
Machine learning provides several paradigms for addressing MDPs in settings where system dynamics are partially known or entirely unknown. Two prominent approaches are reinforcement learning (RL) and approximate dynamic programming (ADP). RL is typically applied in model-free settings, where transition dynamics are unknown, and policies are learned directly through interaction with the environment. Methods such as Q-learning and policy gradient algorithms rely on sampled trajectories to estimate value functions or policy gradients and to iteratively improve decision policies.
Approximate dynamic programming, which has its roots in operations research and control, often assumes access to a simulator or generative model of the system. ADP mitigates the intractability of exact dynamic programming by approximating value functions or policies using function approximation techniques, including linear basis functions, regression trees, and neural networks. These approximations enable iterative solution methods such as fitted value iteration and approximate policy iteration. Although RL and ADP differ in historical origins and terminology, the distinction between the two has become increasingly blurred, particularly with the adoption of deep learning–based methods.
IL provides an alternative paradigm for solving MDPs, especially in settings where specifying a reward function is difficult, misaligned with operational objectives, or undesirable from a modelling perspective. Rather than learning through trial-and-error interaction, IL seeks to learn a policy directly from expert-generated state–action trajectories. This expert-driven formulation is particularly appealing in combinatorial optimisation settings, where high-quality decision policies may be available through optimisation solvers or heuristics, but reward design and exploration are costly.
The effectiveness of IL depends critically on the quality, availability, and structure of the expert demonstrations. While simple approaches such as behavioural cloning treat imitation as a supervised learning problem, more advanced methods incorporate expert interaction or infer latent objectives. The design and use of experts, therefore, play a central role in applying IL to MDPs arising from complex optimisation problems.
@@ 3 Imitation Learning: Models and Challenges
IL [hussein2017imitation] is a specialised paradigm within supervised ML in which an agent learns to perform a task by mimicking behaviour demonstrated by an expert. Unlike classical supervised learning, where data points are typically assumed to be independent and identically distributed, IL commonly arises in sequential decision-making settings. In such problems, observations are temporally correlated and depend on previously executed actions.
The term expert refers to the source of demonstrations used by the learning agent and may correspond to a human decision maker, a heuristic, or an exact optimisation algorithm. Demonstrations are typically provided as trajectories $\tau=\{x_{i},a_{i}^{*}\}_{i=1}^{T}$, where $T$ denotes the trajectory length. An expert dataset is then defined as $\mathcal{D}=\{\tau_{j}\}_{j=1}^{N}$, with each state $x_{i}$ generated under the expert policy $\pi^{*}$ and the corresponding action given by $a_{i}^{*}=\pi^{*}(x_{i})$. In an ML context, actions are also referred to as targets or labels. Standard supervised learning methods can be applied to $\mathcal{D}$ to train a model that predicts expert-like actions from observed states. Here, each state–action pair is an independent training example, effectively flattening trajectories, but it does not require explicitly uncoupling or destroying the underlying trajectory structure.
IL provides a data-driven approach for solving complex SDPs with combinatorial action spaces. However, its effectiveness depends critically on access to high-quality expert demonstrations, which are often expensive to obtain when expert actions are generated by optimisation solvers. In addition, IL must scale to high-dimensional decision spaces and address covariate shift, whereby the learned policy encounters states at deployment that differ from those observed during training, leading to compounding errors. This section introduces two core IL paradigms, discusses these challenges, and motivates the use of approximate or imperfect experts.